find the slope and y intercept of a line whose equation is 6x+5=7

Answer by Alan3354(69443)

find the slope and y intercept of a line whose equation is 6x+5=7
-------
6x+5 = 7
6x = 2
x = 1/3
------
x-int is 1/3, or (1/3,0)
No y-int, it's parallel to the y-axis
Slope is infinite, or undefined.